# Does Generative AI speak Nigerian-Pidgin?: Issues about Representativeness and Bias for Multilingualism in LLMs

### v2

@@ Abstract
Nigeria is a multilingual country with 500+ languages. Naija is a Nigerian-Pidgin spoken by approx. 120M speakers in Nigeria and it is a mixed language (e.g., English, Portuguese, Yoruba, Hausa and Igbo). Although it has mainly been a spoken language until recently, there are now various platforms publishing exclusively in Naija such as Naija Wikipedia. However, it is hard to distinguish by non-native from a larger pidgin languages spoken across West Africa known as West African Pidgin English (WAPE)—which is more simplied and understandable by wider audience in Ghana, Nigeria, and Cameroon. BBC news platform publishes exclusively in WAPE to cater for several countries in West Africa. In our paper, we show through statistical analyses and Machine Translation experiments that these two creole varieties do not represent each other (i.e., there are linguistic differences in word order and vocabulary) and Generative AI operates only based on WAPE. In other words, Naija is under-represented in Generative AI, and it is hard to teach LLMs with few examples.
@@
David Ifeoluwa Adelani Mila - Quebec AI Institute McGill University, Canada CIFAR AI Chair david.adelani@mcgill.ca                        A. Seza Doğruözthanks: Equal contribution LT3, IDLab, Universiteit Gent, Gent, Belgium as.dogruoz@ugent.be
Iyanuoluwa Shode Bloomberg ishode@bloomberg.net                        Anuoluwapo Aremu Lelapa AI
@@ 1 Introduction
Between 16th-19th centuries, there were contacts between Europeans and non-Europeans outside Europe. In West Africa, contacts between English and West African languages leading to simplified and mixed languages with linguistic features from several languages. These new forms of languages were lingua francas (i.e., common or bridge languages) that served for a mutual understanding between speakers of different languages for various purposes (e.g., trade, plantation agriculture, mining) (Mufwene, 2024). "Pidgin" and "creole" are used to refer to these languages. Although there is a lack of agreement about the precise definitions and coverage of these terms, pidgin roughly refers to "speech-forms which do not have native speakers, and are therefore primarily used as a means of communication among people who do not share a common language" Muysken et al. (1995). On the other hand, creoles could be assumed as extended pidgins which are more established and have native speakers especially in urban environments Muysken et al. (1995).
Until recently, West African Pidgin English (WAPE) has been mainly a spoken language with many local varieties (e.g., Nigerian Pidgin, Ghanaian Pidgin English, Cameroonian Pidgin English).
Nigeria is a multilingual country in West Africa African hosting over 500 different languages spoken by approx. 220 million people across 371 ethnic tribes Eberhard et al. (2019). It is the sixth most populous country in the world and Africa’s most populous country (also referred to as “The Giant of Africa” because of its teeming population. English is the official language and acquired mostly through formal education in Nigeria Agbo and Plag (2020). The three major tribes in Nigeria with their respective languages include Hausa (spoken by 63M speakers), Igbo (27M speakers), and Yorùbá (42M speakers). Nigerian Pidgin (Naija) is a mix of English with local languages (e.g., Portuguese, Yorùbá, Igbo, Hausa) Balogun (2013a), Oyebola and Ugwuanyi (2023). Naija is widely spoken (approx. 120M speakers) as a first and second language Adelani (2022) around the Southern part of Nigeria (e.g., Lagos and Niger-Delta) with origins going back to the English-Creole Atlantic Krio language family. It is also adopted as the unifying and unofficial language for communication across ethnically diverse groups. According to some researchers (e.g., Muysken et al. (1995)), Naija has evolved into a creole over time and has now native speakers as well. However, it is still referred to as a pidgin among the locals.
@@
Despite the large number of speakers across West Africa countries, WAPE remained as a spoken language until 2017 when the British Broadcasting Company (BBC) launched a news website (West African Pidgin English). It aims to target the diversity of WAPE speakers across different countries.  [^1] https://www.bbc.com/news/world-africa-40975399 Since 2022, Naija is also accepted as one of the languages on Wikipedia  [^2] https://meta.wikimedia.org/wiki/Requests_for_new_languages/Wikipedia_Nigerian_Pidgin. Although they are mutually understandable, there are linguistic and social differences between the two written varieties. For example, WAPE on BBC resembles English in terms of word order and vocabulary (see example (1)) with a simplified grammar (i.e., lacking auxiliary "were"). However, Naija on Wikipedia has a different word order and vocabulary choice (e.g., ”moto” instead of a "car" and ”wund" instead of ”injured" or ”wounded").
Figure 1: WAPE Locations: West African countries whose official language is English (The Gambia, Sierra Leone, Liberia, Ghana and Nigeria), Cameroon (North-Western and South-Western Anglophone region), and Krios immigrants to Equatorial Guinea. Map obtained from Yakpo (2024).
@@
From the sociolinguistic perspective, WAPE is favored mostly by educated Nigerians (who also speak, read and write in English) whereas Naija is used mostly in everyday life and it is more accessible for larger groups regardless of their background (e.g., educational status) (Akande and Salami, 2012). In terms of availability on the Internet, there is much more data about WAPE than Naija (160K sentences vs. 25K sentences) (Ogueji et al., 2021).
There is a growing need for more research for multilingual and low resource languages Doğruöz and Sitaram (2022); Doğruöz et al. (2021) in Generative AI systems. This need is even enhanced for pidgin and/or creole languages due to their high numbers of speakers but lack of data Lent et al. (2022, 2023). However, there are also unresolved issues about to what extent the available data on the Internet represent the language spoken in real life Doğruöz et al. (2023). It is crucially important that the different genres of the same pidgin and creole languages are also represented in these systems to be inclusive and accessible for all speakers/users with diverse backgrounds (e.g., educated vs. less educated).
In our paper, we address these issues for WAPE and Naija in the Nigerian context with the following contributions. We introduce Warri  [^3] Warri is the name of a city in the South-South region of Nigeria popular for the use of Naija. as a new MT evaluation data set including written WAPE on BBC news website and written Naija on Wikipedia. We also had interviews with the Naija Wikipedia writers to understand their motives and regulations about written Naija.   [^4] We will release the dataset in CC-BY-4.0 license. Our paper is the first to systematically analyze the similarities and differences between WAPE and Naija. Through a Machine Translation experiment, we find that Generative AI models (e.g., GPT-4o and LLaMa 3.1 8B (Touvron et al., 2023)) are biased towards WAPE do not include Naija despite large numbers of speakers. Further analysis shows that LLMs are hard to teach with few examples (e.g., 5-shots) to generate in Naija.
2 Related Work and Links with Our Study
@@
Generative AI systems depend on the availability of large data sets on the Internet. However, this assumption does not consider the representativeness of the variation in the available data sets which is especially difficult for multilingual and low resource languages Doğruöz et al. (2023). While developing language technologies for multilingual and low resource languages, it is crucially important to be aware of the linguistic variation (e.g., WAPE and Naija) within these languages and aim for representing the variation in a balanced way to prevent potential bias.
To investigate potential bias in the Nigerian context, the first step is to establish to what extent WAPE and Naija are similar or different from each other linguistically.
@@ 3 Warri MT benchmark dataset for WAPE and Naija
Table 1: Warri and other datasets: Warri is only used for evaluation in zero or few-shot (e.g. 5) setting. Warri (multi-way) have the same sentences in both WAPE and Naija (PCM, the ISO 639-3 code) pidgins unlike Warri (single-way). We label each dataset based on the specified pidgin assigned by the creators of the dataset.
To establish the Warri data set, we used a portion of WAPE BBC news data, previously used in MasakhaNER dataset (Adelani et al., 2021b)—a named entity recognition dataset with available untokenized texts, and downloaded Naija Wikipedia data from the HuggingFace. [^5] We make use of the 20231101 version, https://huggingface.co/datasets/wikimedia/wikipedia After the data collection, we created the parallel data set in English by recruiting two bilingual speakers. They translated about $505$ sentences from WAPE BBC data and Naija Wikipedia data into English. In this way, we maintained high-quality datasets by preventing the translators from mixing the features of the two pidgins into one. However, this also introduced a new obstacle (i.e., comparison of two test sets from slightly different domains (news vs. Wikipedia).
@@
Table 1 provides the details of our new Warri dataset, containing a single-way parallel sentences translated by two native speakers from WAPE BBC to English, and the multi-way parallel sentences where the English sentences have parallel translation in both WAPE BBC and Naija Wikipedia. Our test set composed of $500$ sentences and the remaining five sentences were for few-shot/in-context learning for LLMs.
@@
    Bible: We found two Naija Bibles online, the first one was translated by Wycliffe Bible Translators, and is part of the freely available eBible corpus (Akerman et al., 2023), and Naija Wikipedia writers also assert that this conforms with Naija rather than WAPE. The other translation was created by the Mercy Christian Ministry International (MCMI),  [^6] https://nigerianpidgin-bible.yolasite.com/ which is written to be closely similar to an African languages in Nigeria including the use of underdot diacritics as they are used in the Igbo language (e.g., “Wọd” for “Word”). Table 2 shows an example of a sentence in the two writing styles which shows there is no standard in the style of writing Naija. We focus on our analysis on Wycliffe since the MCMI does not have the complete bible online. We divided the bible data into 31,051/1,500/1,500 TRAIN/DEV/TEST split.
@@
    UD-Pidgin: This is based on universal dependecy (UD) project for Naija Caron et al. (2019). The data is based on transcript of audio dialog between two Naija speakers.  [^7] https://github.com/UniversalDependencies/UD_Naija-NSC. We divided the data into 6,241/ 1,500/ 1,500 TRAIN/DEV/TEST split.
@@
    MAFAND: This is based on the news domain. The news articles were obtained from English Daily Trust newspaper, and translated to Naija Adelani et al. (2022). We make use of the same split as the MAFAND corpus with 4,790/1,484/1,564. Unlike the other datasets, it can be considered as “general domain” similar to Wikipedia.
Aside from parallel corpora, large amount of WAPE unlabelled texts have been collected in literature  from BBC Ogueji et al. (2021) to train language models such as AfriBERTa. The AfriBERTa corpus more than 160,000 sentences. Other sources of data for Naija are often smaller e.g., Naija tweets Muhammad et al. (2022). However, in our analysis, we primarily focus on parallel data sources.
@@ 4 Experimental setup
We conduct three types of experiments to find out if WAPE and Naija888pcm is the ISO 639-3 code for Naija pidgins are similar to each other: (1) Statistical analysis of the texts obtained from different datasets to find out how similar they are to English and to each other. (2) Cross-corpus zero-shot transfer results when a machine translation (MT) model is trained on one dataset and evaluated on the another. We expect domains that are similar should have a higher performance Adelani et al. (2021a); Lee et al. (2022). Similarly, we expect transfer results to be higher if the creoles are similar in writing. (3) Prompting an LLM to find out whether WAPE or Naija is represented in Generative AI. We compare the result to the evaluation of Warri MT benchmark dataset when trained on Mafand.
@@ 4.1 Statistical analysis of the texts
First, we compute lexical similarity between the English portion of each dataset and Pidgin by measuring Jaccard similarity (in percentage) for each corpus unigram, bigram, and trigram tokens. Secondly, we compute the Levenshtein distance (Levenshtein, 1965) which is an edit distance between each English test sentences and their translations to WAPE and Naija pidgins. Finally, we make use of three additional text generation metrics to measure their similarity to English: BLEU (Papineni et al., 2002), ChrF++ (Popović, 2017) and BERTScore (Zhang et al., 2020). BLEU and ChrF++ are n-gram matching metrics, while BLEU focuses on word-level matching, ChrF++ helps with evaluating character-level differences which are more common for Pidgin, thus more reliable. BERTScore is an embedding-based metric that can measure the semantic relationship between the sentence embeddings of two sentences, thus, they provide better correlation with human judgements.
@@ 4.2 Cross-corpus zero-shot transfer results
Here, we evaluate the performance of training an MT model e.g., on MAFAND, and evaluate on Bible, JW300, UD, BBC, and Wikipedia test sets. We perform evaluation based on ChrF++ due to it’s reliability on capturing the character-level differences between Pidgin and English. Following Adelani et al. (2022), we leveraged a pre-trained model to train an MT model by fine-tuning M2M-100 (418M) on each source data, and evaluated on the remaining test sets of our datasets. In our experiment, only four datasets have training data: Bible, JW300, UD, and MAFAND. The Warri benchmark is only for evaluation.
@@ 5.1 Statistical analysis results
In Table 3, by computing a lexical similarity between the $n$-gram tokens of each genre, we show that WAPE in both news (BBC) and Wikipedia domains consistently have a higher Jaccard similarity score with its parallel English corpus for all $n$-grams, compared to other datasets with the Naija Pidgin label. For example, the unigram similarity score for WAPE was around $0.712-0.802$ while the others are much lower between $0.133$ (Bible) and $0.554$ (Mafand). Interestingly, we find UD, Mafand and Warri Wiki having similar Jaccard similarity.
Furthermore, Levenshtein distance provides an additional evidence of a clear difference between WAPE and Naija. We find that it takes more than twice edit-distance to transform the English sentences to Naija (Wiki) than to WAPE (BBC) and WAPE (Wiki). Naija (Wiki) requires more edits in characters, which shows that it is farther from English compared to the WAPE pidgin. In other words, these two pidgins are quite different than each other linguistically. Similarly, we find longer Levenshtein distance for other datasets: JW300 Mafand, and Bible with $56.5$, $58.0$ and $88.5$ respectively. On the otherhand, UD dataset has a shorter Levenshtein distance compared to others which we attribute to the shorter utterances of the dataset (see Table 1).
Finally, our experiments on text generation metrics such as BLEU, ChrF++ and BERTScore shows that WAPE (BBC) is more similar to English than any of the other Pidgin datasets we evaluated. We find BLEU to be less reliable for this evaluation, achieving only $0.8$ for the Bible while ChrF achieve relatively higher score. We attribute this to several character-level differences between the Bible Pidgin and the English. In general, we find higher score for both WAPE (BBC) and WAPE (Wiki) achieving $65.2-73.4$ ChrF++, while Naija (Wiki) achieved $51.3$. BERTScore evaluation also confirmed this reaching a score of $90.5$ for WAPE while Naija achieved $79.8$.
@@
Table 4: Evaluation on Warri dataset: single-way and multi-way parallel (same sentences translated to both genres) test sets: We compared the performance of MT to different genres using GPT-4-Turbo and adapted M2M-100 (418M) from Mafand training set.

### v4

@@ Abstract
Nigeria is a multilingual country with 500+ languages. Naija is a Nigerian Pidgin spoken by approximately 120M speakers and it is a mixed language (e.g., English, Portuguese, Yoruba, Hausa and Igbo). Although it has mainly been a spoken language until recently, there are some online platforms (e.g., Wikipedia), publishing in written Naija as well. West African Pidgin English (WAPE) is also spoken in Nigeria and it is used by BBC to broadcast news on the internet to a wider audience not only in Nigeria but also in other West African countries (e.g., Cameroon and Ghana). Through statistical analyses and Machine Translation experiments, our paper shows that these two pidgin varieties do not represent each other (i.e., there are linguistic differences in word order and vocabulary) and Generative AI operates only based on WAPE. In other words, Naija is underrepresented in Generative AI, and it is hard to teach LLMs with few examples. In addition to the statistical analyses, we also provide historical information on both pidgins as well as insights from the interviews conducted with volunteer Wikipedia contributors in Naija.
@@
David Ifeoluwa Adelani Mila - Quebec AI Institute McGill University, Canada CIFAR AI Chair david.adelani@mcgill.ca                        A. Seza Doğruözthanks: Equal contribution LT3, IDLab, Universiteit Gent as.dogruoz@ugent.be
Iyanuoluwa Shode Bloomberg ishode@bloomberg.net                        Anuoluwapo Aremu University of Trento Lelapa AI
@@ 1 Introduction
Between 16th-19th centuries, there were contacts between Europeans and non-Europeans outside Europe. In West Africa, contacts between English and West African languages led to simplified and mixed languages combining linguistic features from several languages. These new forms of languages were lingua francas (i.e., common or bridge languages) that served for a mutual understanding between speakers of different languages for various purposes (e.g., trade, plantation agriculture, mining) (Mufwene, 2024). The terms "pidgin" and "creole" are used to refer to these languages. Although there is a lack of agreement about the precise definitions and coverage of these terms, pidgin roughly refers to the "speech-forms which do not have native speakers, and are therefore primarily used as a means of communication among people who do not share a common language" Muysken et al. (1995). Creoles, on the other hand, are assumed to be extended pidgins which are more established and have native speakers especially in urban environments Muysken et al. (1995).
Nigeria is a multilingual country in West Africa hosting over 500 different languages spoken by approximately 220 million people across 371 ethnic tribes Eberhard et al. (2019). It is the sixth most populous country in the world and Africa’s most populous country. English is the official language and acquired mostly through formal education in Nigeria Agbo and Plag (2020). The three major tribes in Nigeria with their respective languages include Hausa (spoken by 63M speakers), Igbo (27M speakers), and Yorùbá (42M speakers). Nigerian Pidgin (Naija) is a mix of English with local languages (e.g., Portuguese, Yorùbá, Igbo, Hausa) Balogun (2013a), Oyebola and Ugwuanyi (2023). Naija is widely spoken (approx. 120M speakers) as a first and second language Adelani (2022) around the Southern part of Nigeria (e.g., Lagos and Niger-Delta) with origins going back to the English-Creole Atlantic Krio language family. It is also adopted as the unifying and unofficial language for communication across ethnically diverse groups. According to some researchers (e.g., Muysken et al. (1995)), Naija has evolved into a creole over time and has now native speakers as well. However, it is still referred to as a pidgin among the locals.
@@
Until recently, West African Pidgin English (WAPE) has been mainly a spoken language with many local varieties (e.g., Nigerian Pidgin, Ghanaian Pidgin English, Cameroonian Pidgin English). Despite the large number of speakers across West Africa countries, WAPE remained as a spoken language until 2017 when the British Broadcasting Company (BBC) launched a news website (West African Pidgin English). It aims to target the diversity of WAPE speakers across different countries.  [^1] https://www.bbc.com/news/world-africa-40975399 Since 2022, Naija is also accepted as one of the languages on Wikipedia  [^2] https://meta.wikimedia.org/wiki/Requests_for_new_languages/Wikipedia_Nigerian_Pidgin. Although they are mutually understandable, there are linguistic and social differences between the two written varieties. For example, (WAPE on BBC website) resembles English in terms of word order and vocabulary (see example (1)) with a simplified grammar, lacking auxiliary "were"). However, Naija on Wikipedia has a different word order and vocabulary choice (e.g., ”moto” instead of a "car" and ”wund" instead of ”injured" or ”wounded").
Figure 1: WAPE Locations: West African countries whose official language is English (The Gambia, Sierra Leone, Liberia, Ghana and Nigeria), Cameroon (North-Western and South-Western Anglophone region), and Krios immigrants to Equatorial Guinea. Map is obtained from Yakpo (2024).
@@
From the sociolinguistic perspective, WAPE is favored mostly by educated Nigerians (who also speak, read and write in English) whereas Naija is used mostly in everyday life and it is more accessible for a larger audience  (Akande and Salami, 2012). In terms of the resources, there is much more data on the Internet about WAPE than Naija (160K sentences vs. 25K sentences) (Ogueji et al., 2021).
There is a growing need for more research for multilingual and low resource languages Doğruöz and Sitaram (2022); Doğruöz et al. (2021) in Generative AI systems. This need is even enhanced for pidgin and/or creole languages due to their high numbers of speakers but lack of data Lent et al. (2022, 2023). However, there are also unresolved issues about to what extent the available data on the Internet represent the language spoken in real life Doğruöz et al. (2023). It is crucially important that the different genres of the same pidgin and creole languages are also represented in these systems to be inclusive and accessible for all speakers/users with diverse backgrounds.
In our paper, we address these issues for WAPE and Naija in the Nigerian context with the following contributions. We introduce Warri  [^3] Warri is the name of a city in Delta State in the South-South region of Nigeria where Naija is widely spoken as a new MT evaluation data set including written WAPE on the BBC website and written Naija on Wikipedia. We also had interviews with the Naija Wikipedia contributors to understand their motives and regulations about written Naija. Our paper is the first to systematically analyze the similarities and differences between WAPE and Naija in Nigeria. Through a Machine Translation (MT) experiment, we find that Generative AI models (e.g., GPT-4o and LLaMa 3.1 8B (Touvron et al., 2023)) are biased towards WAPE and they do not include Naija despite large numbers of speakers. Further analysis shows that LLMs are hard to teach with few examples (e.g., 5-shots) to generate text in Naija. For reproducibility purposes, we release the Warri dataset and our evaluation code on GitHub under CC-BY-4.0 license. [^4] https://github.com/McGill-NLP/Naija-representation-in-LLMs
2 Related Research
@@
Generative AI systems depend on the availability of large data sets on the Internet. However, this assumption does not consider the representativeness of the variation in the available data sets which is especially difficult to obtain for multilingual and low resource languages Doğruöz et al. (2023). While developing language technologies for multilingual and low resource languages, it is crucially important to be aware of the linguistic variation (e.g., WAPE and Naija) within these languages and aim for representing the variation in a balanced way to prevent potential bias.
To investigate the potential bias in the Nigerian context, the first step is to establish to what extent WAPE and Naija are similar or different from each other linguistically.
@@ 3 Warri MT benchmark dataset for WAPE and Naija
Table 1: Warri and other datasets: Warri is only used for evaluation in zero or few-shot (e.g. 5) setting. Warri (multi-way) have the same sentences in both WAPE and Naija (PCM is the ISO 639-3 code) pidgins unlike Warri (single-way). We label each dataset based on the specified pidgin assigned by the creators of the dataset.
To establish the Warri data set, we used a portion of WAPE BBC news data, previously used in MasakhaNER dataset (Adelani et al., 2021b). It is a Named Entity Recognition (NER) dataset with available untokenized texts. We downloaded the Naija Wikipedia data from the Hugging Face. [^5] We make use of the 20231101 version, https://huggingface.co/datasets/wikimedia/wikipedia After the data collection, we created a parallel data set in English by recruiting two bilingual speakers. They translated about $505$ sentences from WAPE BBC data and Naija Wikipedia data into English. In this way, we maintained high-quality datasets by preventing the translators from mixing the features of the two pidgins into one. However, this also introduced a new obstacle (i.e., comparison of two test sets from slightly different domains (news vs. Wikipedia)).
@@
Table 1 provides the details of our new Warri dataset, containing single-way parallel sentences (translated from WAPE BBC to English by two native speakers), and multi-way parallel sentences (where the English sentences have parallel translations in both WAPE BBC and Naija Wikipedia). Our test set composed of $500$ sentences and the remaining five sentences were for few-shot/in-context learning for LLMs.
@@
    Bible: We found two Naija Bibles online. The first one was translated by Wycliffe Bible Translators, and is part of the freely available eBible corpus (Akerman et al., 2023). Naija Wikipedia contributors also agree that this Bible conforms with Naija rather than WAPE. The other Bible translation was created by the Mercy Christian Ministry International (MCMI),  [^6] https://nigerianpidgin-bible.yolasite.com/ which is written to be closely similar to an African languages in Nigeria including the use of underdot diacritics as they are used in the Igbo language (e.g., “Wọd” for “Word”). Table 2 shows the two Naija Bible styles but there is not an established standard. We focus on our analysis on Wycliffe Bible (Naija 1) since the MCMI Bible (Naija 2) does not have the complete Bible online. We divided the Wycliffe Bible data into 31,051/1,500/1,500 TRAIN/DEV/TEST split.
@@
    UD-Pidgin: This is based on the Universal Dependecy (UD) project for Naija Caron et al. (2019). The data is based on the transcript of a conversationbetween two Naija speakers.  [^7] https://github.com/UniversalDependencies/UD_Naija-NSC We divided the data into 6,241/ 1,500/ 1,500 TRAIN/DEV/TEST split.
@@
    MAFAND: This is based on the news domain. The news articles were obtained from English Daily Trust newspaper (published in Nigeria), and translated to Naija Adelani et al. (2022). We make use of the same split as the MAFAND corpus with 4,790/1,484/1,564. Unlike the other datasets, it can be considered as “general domain” similar to Wikipedia.
Aside from parallel corpora, large amounts of WAPE unlabelled texts have been collected in literature  from BBC Ogueji et al. (2021) to train language models such as AfriBERTa. The AfriBERTa corpus has more than 160,000 sentences. Other sources of data for Naija are often smaller (e.g., Naija tweets Muhammad et al. (2022)). However, we primarily focus on the parallel data sources for our analyses.
@@ 4 Experimental setup
We conduct three types of experiments to find out if WAPE and Naija are similar to each other: (1) Statistical analyses of the texts obtained from different datasets to measure their similarity to English and to each other. (2) Cross-corpus zero-shot transfer results when an MT model is trained on one dataset and evaluated on another. We expect domains that are similar should have a higher performance Adelani et al. (2021a); Lee et al. (2022). Similarly, we expect transfer results to be higher if the pidgins are similar in terms of writing. (3) Prompting an LLM to find out whether WAPE or Naija is represented in Generative AI. We compare the results to the evaluation of Warri MT benchmark dataset when trained on Mafand.
@@ 4.1 Statistical analysis of the texts
First, we compute the lexical similarity between the English portion of each dataset and Pidgin by measuring Jaccard similarity (in percentage) for each corpus unigram, bigram, and trigram tokens. Secondly, we compute the Levenshtein distance (Levenshtein, 1965) which is an edit distance between each English test sentences and their translations to WAPE and Naija. Finally, we make use of three additional text generation metrics to measure their similarity to English: BLEU (Papineni et al., 2002), ChrF++ (Popović, 2017) and BERTScore (Zhang et al., 2020). BLEU and ChrF++ are n-gram matching metrics, while BLEU focuses on word-level matching, ChrF++ helps with evaluating character-level differences which are more common for Pidgin. Therefore, it is more reliable. BERTScore is an embedding-based metric that measures the semantic relationship between the sentence embeddings of two sentences. Therefore, it has better correlations with the human judgments.
@@ 4.2 Cross-corpus zero-shot transfer results
We evaluate the performance of training an MT model on a source corpus and evaluate the performance on a target corpus. The source corpus are MAFAND, Bible, JW300, and UD, while the target corpus can be one of source corpora, and the Warri dataset i.e. single-way and multi-way test sets. We perform an evaluation based on ChrF++ due to its reliability about capturing the character-level differences between Pidgin and English. Following Adelani et al. (2022), we leveraged a pre-trained model to train an MT model by fine-tuning M2M-100 (418M) on each source data, and evaluated on the remaining test sets of our datasets.
@@ 5.1 Statistical analysis results
In Table 3, by computing a lexical similarity between the $n$-gram tokens of each genre, we show that WAPE in both news (BBC) and Wikipedia domains consistently have a higher Jaccard similarity score with its parallel English corpus for all $n$-grams, compared to other datasets with the Naija label. For example, the unigram similarity score for WAPE was around $0.712-0.802$ while the others are much lower between $0.133$ (Bible) and $0.554$ (Mafand). UD, Mafand and Warri Wiki data sets have similar Jaccard similarities.
Furthermore, Levenshtein distance provides an additional evidence of a difference between WAPE and Naija. It takes more than twice edit-distance to transform the English sentences to Naija (Wiki) than to WAPE (BBC) and WAPE (Wiki). Naija (Wiki) requires more edits in characters, which shows that it is farther from English compared to the WAPE. In other words, these two pidgins are quite different than each other linguistically. Similarly, we find longer Levenshtein distance for other datasets: JW300 Mafand, and Bible with $56.5$, $58.0$ and $88.5$ respectively. On the otherhand, UD dataset has a shorter Levenshtein distance compared to others which we attribute to the shorter utterances of the dataset (see Table 1).
Finally, our experiments on text generation metrics (e.g., BLEU, ChrF++ and BERTScore) show that WAPE (BBC) is more similar to English than any of the other Pidgin datasets we evaluated. We find BLEU to be less reliable for this evaluation, achieving only $0.8$ for the Bible while ChrF achieve relatively higher scores. We attribute this result to several character-level differences between the Bible Pidgin and the English. In general, we find higher scores for both WAPE (BBC) and WAPE (Wiki) ($65.2-73.4$ ChrF++) than Naija (Wiki) ($51.3$). BERTScore evaluation also confirmed this finding by reaching to a score of $90.5$ for WAPE and $79.8$ for Naija.
@@
Table 4: Evaluation on Warri dataset: single-way and multi-way parallel (same sentences translated to both pidgins) test sets: We compared the performance of MT to different genres using GPT-4-Turbo and adapted M2M-100 (418M) from Mafand training set.